OpenStudy (anonymous):

How do you solve? 3x + 3y + 2z = –5 x + y + 3z = –4 2x + 2y + 5z = –7

OpenStudy (anonymous):

using matrices

OpenStudy (anonymous):

can you set up the augmented matrix?

OpenStudy (unklerhaukus):

add the second and third equations, and find the difference with the first equation, hence solve for z

OpenStudy (anonymous):

or you can do matrices.

OpenStudy (callisto):

3x + 3y + 2z = –5 -(1) x + y + 3z = –4 -(2) 2x + 2y + 5z = –7 -(3) (2)x(-2) + (2) -2 (x + y + 3z) + (2x + 2y + 5z) = -2(-4) + (-7) -z = 1 z= -1 Sub z=-1 into the equations. Solve x and y. If I haven't got that wrong, there should be infinitely many solutions for x and y. Sorry, I made a mistake in the deleted post.